An investor invested $13000 part a 8% and the rest at 12%. Find the amount invested at each rate if the annual income from the two investments is $1240
Answer by mananth(16949)

Part I 8.00% per annum ------------- Amount invested =x
Part II 12.00% per annum ------------ Amount invested = y
13000
Interest----- 1240.00

Part I 8.00% per annum ---x
Part II 12.00% per annum ---y
Total investment
x + 1 y= 13000 -------------1
Interest on both investments
8.00% x + 12.00% y= 1240
Multiply by 100
8 x + 12 y= 124000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -104000.00
Add this to (2)
0 x 4 y= 20000
divide by 4
y = 5000
Part I 8.00% $ 8000
Part II 12.00% $ 5000

CHECK
8000 --------- 8.00% ------- 640.00
5000 ------------- 12.00% ------- 600.00
Total -------------------- 1240.00

m.ananth@hotmail.ca